OpenStudy (anonymous):

write the product or quotient in standard form 1+3i/2i

OpenStudy (anonymous):

1/2(3+2) 5/2

OpenStudy (anonymous):

Multiply NUM and DEN by i = i(1+3i)/2i^2 = i(1+3i)/(-2) = -(i+3i^2)/2 = -(i-3)/2 = (3-i)/2